Determine whether the equation defines y as a function of x.
x+y=9
and
x^2+y^2=1
and
x=y^2

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



In each case, graph the relation. For each graph determine if you can find a vertical line that intersects the graph in two places. If you can, then you cannot call the relation a function. A function must have a unique value for any given input value. It is perfectly ok if two different input values have the same output, but it is not ok if for one given input, the output is "maybe this, maybe that".

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Determine whether the equation defines y as a function of x.
x+y=9
y = -x+9
y is a function of x.
----------

x^2+y^2=1
y = +-sqrt(-x^2+1)
y is not a function as there are two y-value for some values of "x".
-------------------------------
and
x=y^2
y = +-sqrt(x)
===================
Cheers,
Stan H.